OpenStudy (anonymous):

Use the triangle at the right. Find the length of the missing side. 1. a = 16, b = 63 2. b = 2.1, c = 2.9

OpenStudy (anonymous):



OpenStudy (lukebluefive):

Since this is a right triangle, the Pythagorean theorem applies, which states: \[a^{2} + b^{2} = c^{2}\] So, for the first problem: \[c = \sqrt{16^{2} + 63^{2}}\]

OpenStudy (anonymous):

1.79 2.5

OpenStudy (anonymous):

Ok so what'sthe second one?

OpenStudy (anonymous):

how did you get the first one? can you show me?

OpenStudy (anonymous):

10

OpenStudy (lukebluefive):

The second problem is slightly more complicated. Rewrite the equation so that "a" is on the left side by itself: \[a^{2} = c^{2} - b^{2}\] Then take the square root of both sides: \[a = \sqrt{ c^{2} - b^{2}}\] Now insert the values for "b" and "c" that you were given: \[a = \sqrt{ (2.9)^{2} - (2.1)^{2}}\]

OpenStudy (lukebluefive):

I apologize for skipping steps in the first problem. Here are all of the steps for soling the first problem. \[a^{2} + b^{2} = c^{2}\] Moving "c" to the left side of the equation: \[c^{2} = a^{2} + b^{2}\] Squaring both sides: \[c = \sqrt{a^{2} + b^{2}}\] Substituting the given values: \[c = \sqrt{16^{2} + 63^{2}}\] Hope this helps.

OpenStudy (lukebluefive):

Just noticed there is one more step I made implicitly. The following are true: \[\sqrt{a^{2}} = a\] \[\sqrt{c^{2}} = c\]